OpenStudy (anonymous):

will medal and fan...............................................A piece of tile is 4 1/2 inches long and 3 1/4 inches wide. What is the area of the tile?

OpenStudy (anonymous):

14.625

OpenStudy (texaschic101):

area = length x width 3 1/4 * 4 1/2 -- turn to improper fractions 13/4 * 9/2 = 117/8 = 14 5/8

OpenStudy (anonymous):

multiply the two

OpenStudy (anonymous):

k thanks to the both of you